Calculate the future worth of the purchase shown below using an interest rate of 6% per year.

First Cost: -$17,000

Operating Cost: -$3,000/YR

Salvage Value: $2,000

Useful Life: 7 Years
Answer #1

ANSWER:

I = 6%

FW = first cost(f/p,i,n) + operating cost(f/a,i,n) + salvage value

fw = -17,000(f/p,6%,7) - 3,000(f/a,6%,7) + 2,000

fw = -17,000 * 1.504 - 3,000 * 8.394 + 2,000

fw = - 25,568 - 25,182 + 2,000

fw = - 48,750

so the future worth is -$48,750